OpenStudy (anonymous):

Simplify 4√6/√30 by rationalizing the denominator. Please show steps!

OpenStudy (anonymous):

\[\frac{ 4√6 }{ \sqrt{30} }\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

\[\frac{ 4\sqrt6 }{ \sqrt{30} } \times \frac{ \sqrt{30} }{ \sqrt{30} }\] \[\frac{ 4\sqrt6 (\sqrt{30}) }{ \sqrt{30}\times \sqrt{30} }\] \[\frac{ 4\sqrt{180} }{ 30 }\] \[\frac{ 24\sqrt5 }{ 30 }\] \[\frac{ 4\sqrt5 }{ 5 }\]

OpenStudy (anonymous):

@kausarsalley could you help me one more time?

OpenStudy (anonymous):

okay...what is the question??

OpenStudy (anonymous):

\[(2\sqrt{5} + 3\sqrt{7})^2 \]

OpenStudy (anonymous):

okay....the simplest way to expand is to go by this... \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

It also says to justify each step? Does that mean I have to explain each step?

OpenStudy (anonymous):

using that, \[(2\sqrt5+3\sqrt7)^2=(2\sqrt5)^2+2(2\sqrt5)(3\sqrt7)+(3\sqrt7)^2\]

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

its not done with simplifying..

OpenStudy (anonymous):

or can you continue from there??

OpenStudy (anonymous):

Oh, I was just saying.

OpenStudy (anonymous):

I'm not sure what to do.

OpenStudy (anonymous):

\[(2^2(\sqrt5)^2)+4\sqrt5(3\sqrt7)+(3^2(\sqrt7)^2\] \[20 + 12\sqrt{35}+63\] \[83+12\sqrt{35}\]

OpenStudy (anonymous):

do you understand what i have done??

OpenStudy (anonymous):

I think I might, is that the final answer?

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

sorry for the late reply.... yw